OpenStudy (anonymous):

Given the quantities a = 9.6 m, b = 6.5 s, c = 21 m/s, what is the value of the quantity d=a^3/cb^2?

OpenStudy (anonymous):

What are the steps needed to answer this

OpenStudy (anonymous):

set up the equation and insert all numbers (9.6)^3/(21)(6.5)^2=0.9972

OpenStudy (anonymous):

what would be the unit

OpenStudy (anonymous):

Lets just think that we don't need the numbers then it will be easier for me :) a = m, b = s and c = m/s now we replace the in the formula. \[ d = \frac{a^3}{c*b^2} = \frac{m^3}{\frac{m}{s}*s^2} \] \[ d = \frac{m^3}{\frac{m*s^2}{s}} = \frac{m^{\cancel{3}2}}{\frac{\cancel{ms^{\cancel{2}}{\cancel{s}}}}\]

OpenStudy (anonymous):

I can't see if the latex is rendering properly but the end result should be d = m^2/s

OpenStudy (anonymous):

okay, thank you.